# Geometry-Informed Neural Networks

### v3

@@ 3.2 Generating diverse solutions
The last missing piece to training a generative GINN is making $S$ a diverse collection of solutions. In the typical supervised generative modeling setting, the diversity of the generator is inherited from the diversity of the training dataset. The violation of this is studied under phenomena like mode-collapse in GANs [46]. Exploration beyond the training data has been attempted by adding an explicit diversity loss, such as entropy [2], Coulomb repulsion [47], determinantal point processes [48, 49], pixel difference, and structural dissimilarity [43]. We observe that simple generative GINN models are prone to mode-collapse, which we mitigate by adding a diversity constraint.
Many scientific disciplines require to measure the diversities of sets which has resulted in a range of definitions of diversity [50, 51, 52]. Most start with a distance $d:\mathcal{F}^{2}\mapsto[0,\infty)$, which can be transformed into the related dissimilarity. Diversity $\delta:2^{\mathcal{F}}\mapsto[0,\infty)$ is then the collective dissimilarity of a set [51], aggregated in some way. In the following, we describe these two aspects: the distance $d$ and the aggregation into the diversity $\delta$.
@@
Adopting terminology from Enflo [51], we use the minimal aggregation measure:
@@
This choice is motivated by the concavity property, which promotes uniform coverage of the available space, as depicted in Figure 11. Figure 5 illustrates how it counteracts mode-collapse in a geometric problem. However, Equation 2 is well-defined only for finite sets, so, in practice, we apply $\delta$ to a batch of $k$ i.i.d. sampled shapes $S_{k}=\{G(z_{i})|z_{1},...,z_{k}\overset{iid}{\sim}p_{Z}\}$. We leave the consideration of diversity on infinite sets, especially with manifold structure, to future research.
@@
If $f_{j}$ is an SDF then $\int_{\partial\Omega_{i}}f_{j}(x)^{2}\operatorname{d}\!{x}=\int_{\partial\Omega_{i}}\min_{x^{\prime}\in\partial\Omega_{j}}||x-x^{\prime}||_{2}^{2}\operatorname{d}\!{x}$ (analogously for $f_{i}$) and $d$ is closely related to the chamfer discrepancy [53]. We note that $d$ is not a metric distance on functions, but recall that we care about the geometries they represent. Using appropriate boundary samples, one may also directly compute a geometric distance, e.g., any point cloud distance [53]. However, the propagation of the gradients from the geometric boundary to the function requires the consideration of boundary sensitivity [54], which we leave for future work.
In summary, training a GINN corresponds to solving a constrained optimization problem, i.e. improving the expected objective $O(S)$ and feasibility $C_{i}(S)$ w.r.t. to each geometric constraint $i=1..m$ and the diversity constraint $C_{m+1}(S_{k})=\max(\delta(S_{k})-\delta_{\text{min}},0)$. In practice, we convert this into a sequence of unconstrained optimization problems using the augmented Lagrangian method introduced in Section 4.1.
@@ 4 Experiments
We proceed to demonstrate the proposed GINN framework experimentally on a set of validation problems concluding with a realistic 3D engineering design use case. To the best of our knowledge, data-free shape-generative modeling is an unexplored field with no established baselines, problems, and metrics. In addition to the problems defined and solved in Sections 4.2 and 4.3, we define metrics for each constraint in the main problem as detailed in Appendix C.1. We use these to perform quantitative ablation studies in Appendix C.2, reserving the primary text for the main findings. We proceed with an overview of key experimental considerations with more implementation and experiment details available in Appendix A.
@@ 4.1 Experimental details
To solve the aforementioned constrained optimization problems in Eq. 1, we employ the augmented Lagrangian method (ALM). It is well studied in the classical and more recently deep learning literature and balances the feasibility and optimality of the solution by controlling the influence of each constraint while avoiding the ill-conditioning and convergence issues of simpler methods. Specifically, we use an adaptive ALM proposed by Basir and Senocak [55] which uses adaptive penalty parameters $\mu_{i}$ for each constraint to solve problem 1 as the unconstrained optimization problem $\max_{\lambda}\min_{\theta}\mathcal{L}(\theta,\lambda,\mu)$ where
$\mathcal{L}(\theta,\lambda,\mu):=O(S_{k}(\theta))+\sum_{i=1}^{m+1}\lambda_{i}C_{i}(S_{k}(\theta))+\frac{1}{2}\sum_{i=1}^{m+1}\mu_{i}C_{i}^{2}(S_{k}(\theta))\ .$ (4)
The multipliers $\lambda_{i}$ and the penalty parameters $\mu_{i}$ are updated during training according to Equations (18) - (20). Adaptive ALM allows GINNs to handle different constraints without manual hyperparameter tuning for each loss. Appendix D provides a more detailed introduction and motivation for this approach.
@@
describes properties of a shape that are invariant under deformations, such as the number of connected components or holes. Certain materials and objects display specific topological properties [56, 57, 58], e.g., connectedness, which is a basic requirement for the propagation of forces and by extension manufacturability and structural function.
Despite topological properties being discrete-valued, persistent homology (PH) is a tool that allows to formulate a differentiable loss. In brief, it identifies topological features (e.g., connected components) and quantifies their persistence w.r.t. some filtration function. For our implicit shapes, this is the implicit function $f$ itself. Consequentially, the birth and death of each feature can be matched to a pair of critical points of $f$. Their values can then be adjusted to achieve the desired topology.
In practice, we follow the standard procedure of first discretizing the continuous function onto a cubical complex. We additionally filter cells outside the given design space to prevent undesirable connections. Throughout this section, we use a constraint that encourages connected shapes with a minimal number of holes. We detail PH and our approach in Appendix E.
@@
constitutes another computationally non-trivial design requirement that we consider. Many alternative smoothing energies exist, each leading to different surface qualities [59, 60], but a broad class can be written as the surface integral $\int_{\partial\Omega\setminus\mathcal{I}}e(\kappa_{1}(x),\kappa_{2}(x))\operatorname{d}\!{x}$ of some curvature expression $e:\mathbb{R}^{2}\mapsto\mathbb{R}$. The principal curvatures $\kappa_{1}$ and $\kappa_{2}$, same as other differential-geometric quantities, can be computed from $\nabla_{x}f$ and $H_{x}f$ in closed-form [61]. To solve Plateau’s problem, we use the mean curvature $\kappa_{H}:=\frac{\kappa_{1}+\kappa_{2}}{2}$. In the main experiment, we primarily focus on the surface-strain $E:=\kappa_{1}^{2}+\kappa_{2}^{2}$ and a variant thereof $E_{\text{log}}:=\log(1+E)$ to produce visually appealing shapes.
@@
is required to estimate the surface integrals for smoothness and diversity. We first sample points in the envelope and project them onto the surface using Newton iterations. We then repel the points on the surface to achieve a more uniform distribution similar to Yifan et al. [62]. Finally, we exclude points sampled within a small distance to the interface $\mathcal{I}$, as the surface should not change here. We also begin to sample surface points and compute the surface integrals only after a warm-up phase of 500 iterations. In combination, these aspects lead to a lower variance and better convergence.
@@
Across the experiments, we consider several neural field models. We require these to have a well-defined and non-vanishing second derivative w.r.t. the inputs $x$ to compute the surface normals and curvatures. As the simplest models, we use MLPs with a softplus (i.e., differentiable ReLU) activation. These suffice for the simpler validation experiments, but their simplicity bias is limiting for the physics and the 3D experiment. Following the recommendation of Wang et al. [11], we find SIREN suitable for the physics task. For the main task, we employ WIRE [34] – a generalization of SIREN that replaces sine with Gabor wavelet activation functions, which are localized in both the spatial and frequency domains. We show a qualitative comparison in Figure 7. As the neural field conditioning mechanism, we always use input concatenation (see Section 2.3), denoting the latent space dimension as $\dim(z)$.
4.2 Validation problems
Generative PINN solving an under-determined reaction-diffusion problem.
As a primer to geometry tasks, we begin by illustrating solution multiplicity on an under-determined physics system. While most familiar problems in physics are well-defined, cases exist where, e.g., the initial conditions are irrelevant and general PDE solutions are sought, such as in chaotic systems or animations. We first demonstrate how a PINN can also be extended to provide diverse stationary solutions using a system of reaction-diffusion with no initial condition. Figure 4 illustrates the resulting Turing patterns that continuously morph during latent space traversal. For more detail, we refer to Appendix B.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: A generative PINN producing Turing patterns that morph during latent space interpolation. This is a result of searching for diverse solutions to an under-determined Gray-Scott system.
@@ 4.2 Validation problems
Plateau’s problem is to find the surface $M$ with the minimal area given a prescribed boundary $\Gamma$ (a closed curve in $\mathcal{X}\subset\mathbb{R}^{3}$). A minimal surface is known to have zero mean curvature $\kappa_{H}$ everywhere. Minimal surfaces have boundaries and may contain intersections and branch points [63] which cannot be represented implicitly. For simplicity, we select a suitable problem instance, noting that more appropriate geometric representations exist [64, 32]. Altogether, we represent the surface as $M=\partial\Omega\cap\mathcal{X}$ and the two constraints are: $\Gamma\subset M$ and $\kappa_{H}(x)=0\ \forall x\in M$. The result in Figure 5 qualitatively agrees with the known solution.
@@
Although we mainly focus on the implicit representation, the GINN framework extends to other representations, such as explicit, parametric, or discrete shapes. Here, the GINN learns the height function $f:[-1,1]\mapsto\mathbb{R}$ of a mirror with the interface constraint $f(0)=0$ and that all the reflected rays should intersect at the single point $(0,1)$. The result in Figure 5 approximates the known solution: a parabolic mirror. This is a very basic example of caustics, an inverse problem in optics, which we hope inspires future work combining GINNs and the recent developments in neural rendering techniques.
Obstacle to introduce diversity and connectedness.
@@
Consider a 2D rectangular design region $\mathcal{E}$ with a circular obstacle in the middle. The interface $\mathcal{I}$ consists of two vertical line segments and has prescribed outward-facing normals $\bar{n}$. We seek shapes that connect these two interfaces while avoiding the obstacle. Despite this problem admitting infinitely many solutions, the naive application of the generative softplus-MLP leads to mode-collapse. This is mitigated by employing the additional diversity constraint as illustrated in Figure 5.
w/o diversity Refer to caption w/ diversity Refer to caption
Figure 5: GINN solving three validation problems. (a) Plateau’s problem: the unique minimal surface that attaches to the prescribed boundary. (b) Parabolic mirror: the unique surface that collects reflected rays into a single point. (c) Obstacle: connecting the two interfaces within the allowed design region. A superposition of 16 solutions is shown trained with and without a diversity loss, which is required to avoid mode-collapse.
4.3 Engineering design case study
The problem specification
@@ 4.3 Engineering design case study
is based on an engineering design competition hosted by General Electric and GrabCAD [65]. The challenge was to design the lightest possible lifting bracket of a jet engine subject to both physical and geometrical constraints. Here, we focus only on the geometric requirements: the shape must fit in a provided design region $\mathcal{E}$ and attach to five cylindrical interfaces $\mathcal{I}$: a horizontal loading pin and four vertical fixing bolts. Instead of minimizing the volume subject to a linear elasticity PDE constraint, we minimize the surface smoothness $E$ subject to a topological connectedness constraint. Conceptually, this formulation is similar but avoids the need for a PDE solver in the training loop. These requirements are detailed in Table 1 and illustrated in Figs. 1 and 3.
Single solution
to the above problem is included in Figure 3. The trained GINN model represents a smooth, singly connected shape attaching to the interfaces while remaining within the given design space.
Ablations.
In addition to the qualitative presentation, we perform quantitative evaluation and ablations in Appendix C. Table 3 quantifies the impact of the connectedness constraint and the smoothness objective, as well as some experimental decisions, including the WIRE hyperparameters and eikonal regularization.

### v4

@@ 3.2 Generating Diverse Solutions
The last missing piece to training a generative GINN is making $S$ a diverse collection of solutions. In the typical supervised generative modeling setting, the diversity of the generator is inherited from the diversity of the training dataset. The violation of this is studied under phenomena like mode-collapse in GANs (Che et al., 2017). Exploration beyond the training data has been attempted by adding an explicit diversity loss, such as entropy (Noé et al., 2019), Coulomb repulsion (Unterthiner et al., 2018), determinantal point processes (Chen & Ahmed, 2020; Heyrani Nobari et al., 2021), pixel difference, and structural dissimilarity (Jang et al., 2022). We observe that simple generative GINN models are prone to mode-collapse, which we mitigate by adding a diversity constraint.
Many scientific disciplines require to measure the diversities of sets which has resulted in a range of definitions of diversity (Parreño et al., 2021; Enflo, 2022; Leinster & Cobbold, 2012). Most start with a distance $d:\mathcal{F}^{2}\mapsto[0,\infty)$, which can be transformed into the related dissimilarity. Diversity $\delta:2^{\mathcal{F}}\mapsto[0,\infty)$ is then the collective dissimilarity of a set (Enflo, 2022), aggregated in some way. In the following, we describe these two aspects: the distance $d$ and the aggregation into the diversity $\delta$.
@@
Adopting terminology from Enflo (2022), we use the minimal aggregation measure:
@@
This choice is motivated by the concavity property, which promotes uniform coverage of the available space, as depicted in Figure 13. Figure 5(c) illustrates how it counteracts mode-collapse in a geometric problem. However, Equation 2 is well-defined only for finite sets, so, in practice, we apply $\delta$ to a batch of $k$ i.i.d. sampled shapes $S_{k}=\{G(z_{i})|z_{1},...,z_{k}\overset{iid}{\sim}p_{Z}\}$. We leave the consideration of diversity on infinite sets, especially with manifold structure, to future research.
@@
If $f_{j}$ (analogously $f_{i}$) is an SDF then $\int_{\partial\Omega_{i}}f_{j}(x)^{2}\operatorname{d}\!{x}=\int_{\partial\Omega_{i}}\min_{x^{\prime}\in\partial\Omega_{j}}||x-x^{\prime}||_{2}^{2}\operatorname{d}\!{x}$ and $d$ is closely related to the chamfer discrepancy (Nguyen et al., 2021). We note that $d$ is not a metric distance on functions, but recall that we care about the geometries they represent. Using appropriate boundary samples, one may also directly compute a geometric distance, e.g., any point cloud distance (Nguyen et al., 2021).
In summary, training a GINN corresponds to solving a constrained optimization problem, i.e. improving the expected objective $O(S)$ and feasibility $C_{i}(S)$ w.r.t. to each geometric constraint $i=1..m$ and the diversity constraint $C_{m+1}(S_{k})=\max(\delta(S_{k})-\bar{\delta},0)$. In practice, we convert this into a sequence of unconstrained optimization problems using the augmented Lagrangian method introduced in Section 4.1.
@@ 4 Experiments
We demonstrate the proposed GINN framework experimentally on a range of problems spanning physics, geometry, and engineering, summarized in a problem-constraint matrix (Table 2). To the best of our knowledge, data-free shape-generative modeling is an unexplored field with no established baselines, problems, and metrics. Thus, in addition to the problems defined and solved in Section 4.2, we define metrics for each constraint as detailed in Appendix C.1. We use these to perform quantitative ablation studies in Appendix C.2 and evaluate baselines in Appendix C.4, reserving the primary text for the main findings. We proceed with an overview of key experimental considerations, with more experimental and implementation details available in Appendix B.
@@ 4.1 Experimental Details
To solve the aforementioned constrained optimization problem in Equation 1, we employ the augmented Lagrangian method (ALM). It is well studied in the classical and more recently deep learning literature. ALM balances the feasibility and optimality of the solution by controlling the influence of each constraint while avoiding the ill-conditioning and convergence issues of simpler methods. Specifically, we use an adaptive ALM (Basir & Senocak, 2023) that uses adaptive penalty parameters $\mu_{i}$ for each constraint to solve Equation 1 as the unconstrained optimization problem $\max_{\lambda}\min_{\theta}\mathcal{L}(\theta,\lambda,\mu)$ where
$\mathcal{L}(\theta,\lambda,\mu):=O(S_{k}(\theta))+\sum_{i=1}^{m+1}\lambda_{i}C_{i}(S_{k}(\theta))\\ +\frac{1}{2}\sum_{i=1}^{m+1}\mu_{i}C_{i}^{2}(S_{k}(\theta))\ .$ (4)
The multipliers $\lambda_{i}$ and the penalty parameters $\mu_{i}$ are updated during training according to Equations (18) - (20). Adaptive ALM allows GINNs to handle different constraints without manual hyperparameter tuning for each loss. However, ALM works best if the losses are already on a similar scale. Appendix D provides a more detailed introduction and motivation for this approach.
@@
describes properties of a shape that are invariant under deformations, such as the number of holes. Certain materials and objects display specific topological properties (Moore, 2010; Caplan et al., 2018; Bendsoe & Sigmund, 2011), e.g., connectedness, which is a basic requirement for the propagation of forces and, by extension, manufacturability and structural function.
Despite topological properties being discrete-valued, persistent homology (PH) allows to formulate a differentiable loss. It identifies topological features (e.g., connected components) and quantifies their persistence w.r.t. some filtration function. For our implicit shapes, this is the implicit function $f$ itself. Consequently, the birth and death of each feature can be matched to a pair of critical points of $f$. Their values can then be adjusted to achieve the desired topology.
In practice, we follow the standard procedure of first discretizing the continuous function onto a cubical complex. We filter cells outside the design region $\mathcal{E}$ to prevent invalid connections. In some experiments, we use an additional constraint that minimizes the number of holes. We detail PH and our approach in Appendix E.
@@
is another computationally non-trivial design requirement that we consider. Many alternative smoothing energies exist, each leading to different surface qualities (Westgaard & Nowacki, 2001; Song, 2021), but a broad class of smoothing energies can be written as the surface integral $\int_{\partial\Omega\setminus\mathcal{I}}e(\kappa_{1}(x),\kappa_{2}(x))\operatorname{d}\!{x}$ of some curvature expression $e:\mathbb{R}^{2}\mapsto\mathbb{R}$. The principal curvatures $\kappa_{1}$ and $\kappa_{2}$, same as other differential-geometric quantities, can be computed from $\nabla_{x}f$ and $H_{x}f$ in closed-form (Goldman, 2005). To solve Plateau’s problem, we use the mean curvature $\kappa_{H}:=\frac{\kappa_{1}+\kappa_{2}}{2}$. In the bracket experiment, we use the surface-strain $E:=\kappa_{1}^{2}+\kappa_{2}^{2}$ and a variant thereof $E_{\text{log}}:=\log(1+E)$.
@@
is required to estimate the surface integrals for smoothness and diversity. We describe our sampling strategy in Appendix B, noting that it gives a lower variance of the losses and a better convergence compared to a naive strategy.
@@
Across the experiments, we consider several NF models that primarily differ in their activation function, including softplus, SIREN (Sitzmann et al., 2020), and WIRE (Saragadam et al., 2023). We require all models to have well-defined and non-vanishing first and second derivatives $\nabla_{x}f$ and $H_{x}f$ to compute the iso-level normals and curvatures. As the NF conditioning mechanism, we always use input concatenation (see Section 2.3), denoting the latent space dimension as $\dim(z)$. We continue the model choice discussion in Appendix B.
4.2 Problems
Generative PINN solving an under-determined PDE to demonstrate the generality of the approach.
Although we focus mainly on geometric tasks, we show that the idea of a diversity-constrained data-free neural field generator generalizes to other areas, such as physics. While most familiar problems in physics are well-defined, cases exist where, e.g., the initial conditions are irrelevant and general PDE solutions are sought, such as in chaotic systems or animations. Here, we seek stationary solutions to a reaction-diffusion system with no initial condition. Note that this system has infinitely many solutions. Figure 4 illustrates the resulting Turing patterns that continuously morph during latent space traversal of the trained generative PINN. Appendix B.1 describes the problem in greater detail.
Figure 4: A generative PINN producing Turing patterns that morph during latent space interpolation. This is a result of searching for diverse solutions to an under-determined PDE.
(a) Plateau’s problem
(b) Parabolic mirror
(c) Obstacle
(d) Wheels
Figure 5: GINN applied to four different problems. (a) Plateau’s problem: the unique minimal surface that attaches to the prescribed boundary. (b) Parabolic mirror: the unique curve that collects reflected rays into a single point. (c) Obstacle: connecting the two interfaces within the allowed design region. A superposition of 16 solutions is shown trained with (bottom) and without (top) a diversity loss, which is required to avoid mode-collapse. (d) Wheel: The diverse and structured latent space of a GINN trained to discover cyclically symmetric wheel designs. A larger version is available in Figure 7.
@@ 4.2 Problems
Plateau’s problem is to find the surface $M$ with the minimal area given a prescribed boundary $\Gamma$ (a closed curve in $\mathcal{X}\subset\mathbb{R}^{3}$). A minimal surface is known to have zero mean curvature $\kappa_{H}$ everywhere. Minimal surfaces have boundaries and may contain intersections and branch points (Douglas, 1931) which cannot be represented implicitly. For simplicity, we select a suitable problem instance, noting that more appropriate geometric representations exist (Wang & Chern, 2021; Palmer et al., 2022). Altogether, we represent the surface as $M=\partial\Omega\cap\mathcal{X}$ and the two constraints are $\Gamma\subset M$ and $\kappa_{H}(x)=0\ \forall x\in M$. The result in Figure 5(a) approximates the known solution.
@@
Although we mainly focus on the implicit representation, the GINN framework extends to other representations, such as explicit, parametric, or discrete shapes. Here, the GINN learns the explicit height function $f:[-1,1]\mapsto\mathbb{R}$ of a mirror with the interface constraint $f(0)=0$ and that all the reflected rays should intersect at the focal point $(0,1)$. The result in Figure 5(b) approximates the known solution: a parabolic mirror. This is a very basic example of caustics, an inverse problem in optics, which we hope inspires future work leveraging the recent advancements in neural rendering techniques.
Obstacle to introduce connectedness and diversity constraints.
@@
Consider a 2D rectangular design region $\mathcal{E}$ with a circular obstacle in the middle. The interface $\mathcal{I}$ consists of two vertical line segments and has prescribed outward-facing normals $\bar{n}$. We seek shapes that connect these two interfaces while avoiding the obstacle. Despite this problem admitting infinitely many solutions, the naive application of the generative softplus-MLP with $\dim(z)=1$ leads to mode-collapse. This is mitigated by employing the additional diversity constraint as illustrated in Figure 5(c).
Wheels as a simple design problem.
Consider the domain $X=[-1,1]^{2}$ containing the ring-shaped design region =$\mathcal{E}=\{x\in X|0.2^{2}\leq x_{1}^{2}+x_{2}^{2}\leq 0.8^{2}\}$ and the interface $\mathcal{I}=\partial\mathcal{E}$. As before, the shapes must also satisfy the connectedness and diversity constraints. Additionally, a 5-fold cyclic symmetry constraint is required. We implement this as a soft constraint by sampling a point, rotating it by $\frac{2}{5}\pi$ four times, and minimizing the variance of the implicit function $f$ at these five points. Alternatively, exact symmetry can be imposed using a periodic encoding of the input. The result of traversing the 2D latent space of discovered shapes is shown in Figure 5(d) (larger version in Figure 7).
Jet engine bracket to demonstrate scaling to 3D engineering design problems.
@@
The problem specification is based on an engineering design competition hosted by General Electric and GrabCAD (Kiis et al., 2013). The challenge was to design the lightest possible lifting bracket of a jet engine subject to both physical and geometrical constraints. Here, we focus only on the geometric requirements: the shape must fit in the given freeform design region $\mathcal{E}$ and attach to five cylindrical interfaces $\mathcal{I}$: a horizontal loading pin and four vertical fixing bolts (see the sketch in Figure 1). Instead of minimizing the volume subject to a linear elasticity PDE constraint, we minimize the surface smoothness $E$ subject to a topological connectedness constraint. Conceptually, this formulation is similar but avoids a PDE solver in the training loop, which we address in a follow-up work. These requirements and several GINN solutions are illustrated in Figures 1 and 3, which also explore additional requirements, including the modified surface energy $E_{\text{log}}$, number of holes, and diversity. The result of latent space traversal of the GINN model trained with diversity is illustrated in Figure 6. In all cases, GINNs produce smooth, singly connected shapes that attach to the interfaces while remaining within the given design space. These properties are quantified in Appendix C.
4.3 Discussion
Solution diversity, generalization, and latent structure.
With the diversity constraint, GINNs not only produce multiple solutions, but we also observe the emergence of a latent space structure. This is best seen in Figure 6 using a 2D latent space from which $k=9$ random samples are drawn every training iteration. Traversing the latent space of the trained GINN produces continuously morphing feasible shapes, i.e., the model generalizes. Furthermore, the latent space is organized – the solutions vary consistently over large latent distances, and latent directions account for different aspects. We see similar behavior in all four experiments that include the diversity constraint. However, we find that the learned structure depends on the exact setup of the diversity constraint. In particular, we observe a more pronounced organization emerge for larger $\bar{\delta}$, however, over-specifying it impedes convergence.